OpenStudy (anonymous):

Which of the following equations describes the line shown below? Check all that apply.

OpenStudy (anonymous):

A. y-2=-6/5(x+6) B. y-2=-5/6(x+6) C. y+4=-6/5(x+1) D. y+4=-5/6(x+1) E. y-4=-5/6(x-1) F. y-6=-6/5(x+2)

OpenStudy (anonymous):

@jcpd910 help me this eric1212

OpenStudy (anonymous):

You're not allowed to have two accounts... I'm not going to help this account.

OpenStudy (anonymous):

my other 1 got deleted my computer reset